OpenStudy (anonymous):

The quotient of a number and 2 is the same as the difference of the number doubled and 3. What is the number?

OpenStudy (anonymous):

My bad. LOL. I wrote the right thing. accidentally deleted it.

OpenStudy (anonymous):

Let x be the number. \[{x \over 2}=2x-3\] solve for x.

OpenStudy (anonymous):

how would I solve for x? would I do 2x*2?

OpenStudy (anonymous):

hmm...

OpenStudy (anonymous):

Yes, you do 2x*2. I'm sorry. I'm really tired now. Way past bedtime.

OpenStudy (anonymous):

\[{x \over 2} \times 2= 2(2x-3)\]

OpenStudy (anonymous):

that's ok, I hope I'm not a pain. I just don't fully understand it thats all

OpenStudy (anonymous):

No worries. I enjoy explaining.

OpenStudy (anonymous):

Which part do you not understand? The question or how to solve for x?

OpenStudy (anonymous):

2x=4x-6?

OpenStudy (anonymous):

almost correct. except that there is no 2x.

OpenStudy (anonymous):

oh ooops, 4x=4x-6?

OpenStudy (anonymous):

For example, when you multiply 1/2 by 2, you get 1. (because the 2s cancel out each other) In this case, x/2 is multiplied by 2, you get x.

OpenStudy (anonymous):

So would that make it 4 1/2?

OpenStudy (anonymous):

May I ask which syllabus are you learning?

OpenStudy (anonymous):

you are*

OpenStudy (anonymous):

I do school online, so I get random things thrown at me to figure out

OpenStudy (anonymous):

I totally understand. I'm online schooled too. ;)

OpenStudy (anonymous):

So I'm guessing, American syllabus then? (Cause american and british syllabus have different ways of teaching algebra.)

OpenStudy (anonymous):

Oh...yes American Syllabus

OpenStudy (anonymous):

Alright then. This is the equation that your question is asking you to solve: \[{x \over 2}=2x-3\] First, you must get rid of the 2 below the x (on the right side of the equation.)

OpenStudy (anonymous):

To do that, you multiply the number by itself because (as an example) 1/2 multiplied by 2 is somewhat like 2 divided by 2. When you divide 2 by 2, you get 1. Still following?

OpenStudy (anonymous):

Yes I am.

OpenStudy (anonymous):

\[{1 \over 2} \times 2={2 \over 2}=1\]

OpenStudy (anonymous):

okay. so in this case, we have x, which has an invisible coefficient of 1, so you treat it like a 1.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so instead of writing 1, you write x instead. Like in this case, \[{x \over 2} \times 2 ={2x \over 2}=x\]

OpenStudy (anonymous):

anyways, when you multiply the 2 on one side, you have to multiply the 2 on the other side too, that's how you get... \[x=4x-6\] still following?

OpenStudy (anonymous):

Yes (:

OpenStudy (anonymous):

okay, now you want to isolate x to one side so you minus 4x from both sides. \[x-4x=4x-4x-6\] What do you get now?

OpenStudy (anonymous):

3x=6?

OpenStudy (anonymous):

correct! (you're one step ahead, but that's good so, good job!) Now, you want to find x so you divide both sides by 3. \[{3x \over 3}={6 \over 3}\] can you solve this one? :)

OpenStudy (anonymous):

x/2 ?

OpenStudy (anonymous):

Do one side at a time.

OpenStudy (anonymous):

\[{3x \over 3}=?\]

OpenStudy (anonymous):

oh... 3/3=0 so x. and 6/3=2 so 2? x=2?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Just to correct you. 3/3 is actually 1.

OpenStudy (anonymous):

but your answer is correct anyway. :)

OpenStudy (anonymous):

Yay!! Thank You so much for the help!!!

OpenStudy (anonymous):

You're welcome! Just keep doing practices and you'll get the hang of it in no time!